The concentration of HCl can be determined by titration with a primary standard such as Na2CO3 (sodium carbonate). The titration reaction is
CO3^(2-) + 2H+ yields H2O + CO2
What is the concentration of HCl if 25.9 mL of HCl was needed to to react with 0.311 g of Na2CO3 to reach the end point?
Answer #1

\(\mathrm{Na}_{2} \mathrm{CO} 3+2 \mathrm{HCl}=2 \mathrm{NaCl}+\mathrm{H} 2 \mathrm{O}+\mathrm{CO} 2\)

Moles of \(\mathrm{Na} 2 \mathrm{CO} 3=\) mass \(/\) molar mass of \(\mathrm{Na} 2 \mathrm{CO} 3\)

\(=0.311 / 105.99=0.002934 \mathrm{~mol}\)

Moles of \(\mathrm{HCl}=2 \times\) moles of \(\mathrm{Na} 2 \mathrm{CO} 3\)

\(=2 \times 0.002934=0.005868 \mathrm{~mol}\)

Concentration of \(\mathrm{HCl}=\) moles \(/\) volume of \(\mathrm{HCl}\)

\(=0.005868 / 0.0259\)

\(=0.227 \mathrm{M}\)